# Irredundant k–Fold Cross–Validation

### v1

@@ Abstract
In traditional k–fold cross–validation, each instance is used ($k\!-\!1$) times for training and once for testing, leading to redundancy that lets many instances disproportionately influence the learning phase. We introduce Irredundant $k$–fold cross–validation, a novel method that guarantees each instance is used exactly once for training and once for testing across the entire validation procedure. This approach ensures a more balanced utilization of the dataset, mitigates overfitting due to instance repetition, and enables sharper distinctions in comparative model analysis. The method preserves stratification and remains model–agnostic, i.e., compatible with any classifier. Experimental results demonstrate that it delivers consistent performance estimates across diverse datasets –comparable to $k$–fold cross–validation– while providing less optimistic variance estimates because training partitions are non–overlapping, and significantly reducing the overall computational cost.
@@ 1 Introduction
The hold–out method [7] improves on this by partitioning the data into two disjoint subsets: one for training and another for testing (typically, 2/32/32 / 3 for training and 1/31/31 / 3 for testing). The hold–out error estimator is unbiased for the generalization error provided that the test data are truly independent of the training data and have not been reused in any form [8, 9]. While computationally inexpensive, the hold–out method is sensitive to how the data are split, resulting in high variance in the estimated metrics.
To address the instability of hold–out, $k$–fold cross–validation was proposed [10]. In this method, the data is partitioned into $k$ disjoint subsets (folds). Each fold serves once as a test set, while the remaining $k-1$ folds are used for training. The final performance estimate is the average across all $k$ iterations. Cross–validation reduces variance and produces more robust estimates, particularly in small datasets. However, its variance is significantly greater than that of resubstitution [11]. As noted by Nadeau & Bengio [12], the use of cross–validation estimators for model selection sparked a debate [13, 14] related to the “no free lunch theorem” [15], highlighting that while cross–validation often works well in practice, it is not guaranteed to perform uniformly well across all learning problems.
[…]
A limiting case of $k$–fold cross–validation is leave–one–out cross–validation [17, 18, 19], where each individual instance serves once as the test instance while the rest are used for training. Although this method yields minimal bias, it can be computationally intensive and prone to high variance, particularly for models with high complexity.
[…]
@@
Each validation strategy entails trade–offs in terms of computational cost, variance, and bias. Hold–out is fast but unstable; leave–one–out is unbiased but often computationally prohibitive; $k$–fold cross–validation balances variance and bias reasonably well but can still be expensive; and bootstrap methods introduce an additional stochastic layer, sometimes at the cost of increased bias, that may not always reflect real–world generalization.
As models increase in complexity and datasets become larger and more heterogeneous, more sophisticated and computationally efficient validation schemes are needed. This paper introduces irredundant $k$–fold cross–validation, a novel approach that aims to combine statistical rigor with computational efficiency. Unlike standard $k$–fold cross–validation, where each sample is used $k\!-\!1$ times for training, irredundant $k$–fold ensures that each sample is used exactly once for training and exactly once for testing. Experimental results demonstrate that irredundant $k$–fold achieves significant computational savings while preserving the quality of performance estimates.
@@
The remainder of this paper is organized as follows: Section 2 describes the irredundant $k$–fold cross–validation in detail and includes, in Section 3, a brief analysis of its bias and variance; Section 4 presents empirical results across several dataset from the UCI repository. Finally, 5 summarizes the main conclusions and outlines future work.
@@ 2.1 Overview
The fundamental difference between standard $k$–fold cross–validation ($\mathsf{kF}$) and the proposed irredundant $k$–fold cross–validation ($\mathsf{IkF}$) lies how many times each instance is used for training. In traditional $\mathsf{kF}$, each instance appears in the training set ($k-1$) out of the $k$ iterations. Therefore, per iteration:
@@
To achieve this, $\mathsf{IkF}$ introduces a subfolding strategy. Each outer fold is divided into $k-1$ subfolds. For each test fold $\mathcal{F}^{i}$, the training set $\mathcal{E}^{i}$ is constructed by selecting one unique, unused subfold from each fold $\mathcal{F}_{j}$ with $j\neq i$. The subfolds may be selected randomly or via heuristic criteria, as long as each is used exactly once across all folds.
@@
$\mathcal{E}^{i}\cap\mathcal{E}^{j}=\emptyset\qquad\mathcal{T}^{i}\cap\mathcal{T}^{j}=\emptyset\qquad\text{for all }i\neq j$ (3)
@@
The proposed $\mathsf{IkF}$ guarantees that each instance appears exactly once in a test set $\mathcal{T}^{i}$ and exactly once in a training set $\mathcal{E}^{i}$. Furthermore, each fold $\mathcal{F}_{i}$ is used exactly once for testing, and each training set $\mathcal{E}^{i}$ is composed of $k\!-\!1$ distinct subfolds, one from each of the other folds. Thus, the size of each training set remains approximately $n/k$.
@@ 2.2 Algorithm Description
The full $\mathsf{IkF}$ procedure is summarized in Algorithm 1. Structurally, it resembles classical $\mathsf{kF}$, but differs in Line #7, where the training set $\mathcal{E}^{i}$ is built by assembling one unused subfold from each of the $k-1$ non–test folds.
@@
3:  Split $\mathcal{F}_{i}$ into $k$--111 subfolds $\mathcal{F}_{i}^{j}$: $\mathcal{F}_{i}=\bigcup_{j=1}^{k-1}\mathcal{F}_{i}^{j}$
@@
6:  $\mathcal{T}^{i}\!\leftarrow\!\mathcal{F}_{i}$triangleright Test set
7:  $\mathcal{E}^{i}\!\leftarrow\!\bigcup_{j\neq i}\mathcal{F}_{j}^{\alpha_{ij}}$ where $\alpha_{ij}\!\in\!\{1,\!\dots\!,k\!-\!1\}$triangleright Training set
@@
9:  $\theta^{i}$leftarrow Evaluate $\mathcal{M}^{i}$ on $\mathcal{T}^{i}$ with $\delta$triangleright Performance
@@
Because each fold is split into $k\!-\!1$ subfolds, the expected size of a subfold is $n/[k(k-1)]$. Hence the irredundant scheme is feasible only if $n\geq k(k-1)$, which is easily met by most real–world data.
@@
Figure 1: Graphical comparison between $k$–fold cross–validation and Irredundant $k$–fold cross–validation.
Figure 1 illustrates the full $\mathsf{IkF}$ construction for $k=3$. To the left, the traditional $k$–fold cross–validation; in the center, the decomposition for irredundant $k$–fold cross–validation into $k$ folds, which of them is again decomposed into $k\!-\!1$ subfolds.
@@ 2.3 Computational cost
[…]
@@
The relative cost reduction depends on the model’s complexity. For linear–time learners (e.g., naïve Bayes), $\mathcal{C}_{\mathsf{kF}}/\mathcal{C}_{\mathsf{IkF}}\!=\!k\!-\!1$. In a typical 10–fold setup, this means up to 9×9$\times$9 × speed–up with respect to the classical $\mathsf{kF}$ approach.
@@ 3 Analysis
In a generic $k$–fold cross-validation scheme—either standard or the proposed irredundant—the dataset $\mathcal{D}$ is partitioned into $k$ mutually–exclusive training–test pairs $(\mathcal{E}_{1},\mathcal{T}_{1}),\;(\mathcal{E}_{2},\mathcal{T}_{2}),\;\dots,\;(\mathcal{E}_{k},\mathcal{T}_{k})$, where each $\mathcal{T}_{i}$ serves as the validation set for the model trained on $\mathcal{E}_{i}$. The fold–wise error estimate and the resulting cross–validation estimator are defined as follows:
@@
In practice, $\widehat{\theta}$ is widely used as a proxy for the generalization error of $h_{\mathcal{D}}$, although this estimator is only unbiased for the expected risk across models trained on subsets of size $|\mathcal{E}_{i}|$, not necessarily for the risk of the final model trained on $\mathcal{D}$ itself. Nonetheless, under common assumptions—such as model stability or large $n$$\widehat{\theta}$ typically provides a reliable and consistent estimate in most applications.
Under mild regularity assumptions, both standard and irredundant $k$–fold cross–validation yield consistent estimators of the generalization risk. Let $R(h_{n})$ denote the risk of a model trained on $n$ i.i.d. samples. If the learning algorithm $\mathcal{M}$ is universally consistent, i.e., $R(h_{n})\to R^{*}$ as $n\to\infty$ for all $\mathbb{P}_{XY}$, and the loss function $\ell$ is bounded, then the cross–validation estimate $\widehat{\theta}$ converges in probability to $R^{*}$. In particular, for the proposed irredundant procedure $\mathsf{IkF}$, the training set size per fold is $n/k$, which grows with $n$ when $k$ is fixed. Therefore, $\mathsf{IkF}$ remains consistent, since each model in the ensemble is trained on increasingly large samples. Moreover, the absence of training–set overlap in $\mathsf{IkF}$ reduces inter–fold dependence, simplifying variance analysis and enhancing reliability in finite samples. Consequently, $\mathsf{IkF}$ inherits the desirable asymptotic properties of classical cross–validation while providing practical advantages in terms of computational efficiency and variance reduction.
@@ 3.1 Bias Analysis
where the expectation $\mathbb{E}[\cdot]$ is taken over the randomness induced by the partitioning of the data. This $\operatorname{Bias}(\hat{\theta})$ quantifies the systematic error introduced by evaluating the model on limited training subsets.
@@ 3.2 Variance Analysis
where the expectation $\mathbb{E}[\cdot]$ is taken over all possible random partitions of the data into training and testing subsets, according to the cross-validation scheme. The quantity $\text{Var}(\hat{\theta})$ quantifies the sensitivity of the estimate to the specific way the data is split into folds.
High variance indicates that the model’s performance estimate is highly sensitive to the chosen data partition, which undermines the robustness of conclusions. In traditional $\mathsf{kF}$ cross–validation, each instance is used for training in $k\!-\!1$ folds and for testing in one fold. Because training sets might be large and largely overlapping, the variance between fold estimates is often low—primarily due to the shared training samples across iterations. However, this overlap reduces the diversity of the training sets and may lead to over–optimistic estimates of generalization.
In general, the use of standard $\mathsf{kF}$ can lead to uneven model behavior, mainly due to redundant learning: since each instance appears in the training set of $k\!-\!1$ folds, the same examples are learned repeatedly. This redundancy artificially lowers the observed variance in performance estimates, as the model benefits multiple times from having seen the same samples. As a result, $\mathsf{kF}$ can underestimate the variance that would be observed in real deployment on truly unseen data [26].
@@
[…]
@@ 3.3 Bias–Variance Trade–off
where the first term, $(\mathbb{E}[\hat{\theta}]-\theta)^{2}$, corresponds to the squared bias of the estimator, and the second term, $\text{Var}(\hat{\theta})$, represents its variance. That is:
$\text{MSE}(\hat{\theta})=\text{Bias}(\hat{\theta})^{2}+\text{Var}(\hat{\theta})$ (12)
@@
Generally, the variance of $\hat{\theta}$ tends to increase with the number of folds $k$—since each test set becomes smaller and each training set slightly more variable—and decrease as $k$ decreases (i.e., when test sets are larger and folds are more stable). At small $k$ (e.g., $k=5$), bias increases because each model is trained on significantly less data, but variance decreases due to lower instability. At the other extreme, leave-one-out cross-validation ($k=N$) yields minimal bias—since each model is trained on nearly all data—but high variance due to high sensitivity to the specific omitted sample.
@@ 4 Experimental Analysis
Table 1: Performance of the Irredundant $k$–fold CV ($\mathsf{IkF}$) and the standard $k$–fold CV ($\mathsf{kF}$) methods across multiple datasets using the Random Forest classifier. #s, #v and #c stand for the number of samples, variables and classes, respectively. For each method, accuracy (Acc), F–score (Fsc), and running Time (in sec.) are reported. Relative ratios $\mathsf{kF}$/$\mathsf{IkF}$ are shown in the last three columns. Last row shows the mean of each column.
Irredundant $k$–Fold Standard $k$–Fold Ratios $\mathsf{kF}/\mathsf{IkF}$
@@
We benchmarked $\mathsf{IkF}$ and $\mathsf{kF}$ on ten publicly available datasets from the UCI Machine–Learning Repository [27]. Several properties make these datasets particularly suitable for the present analysis. First, they vary substantially in sample size, from relatively small datasets (e.g., hepatitis_c with 1,385 samples) to larger datasets (e.g., magic_gamma with 19,020 samples), allowing the evaluation of validation methods across different sizes. Second, the number of predictive variables covers a wide spectrum, from low–dimensional problems (e.g., data_banknote with 4 variables) to high–dimensional ones (e.g., musk_v2 with 166 variables), providing a testbed for studying the impact of dimensionality. Third, the classification tasks include both binary and multiclass problems, ensuring that the methods are assessed under varying levels of label complexity. These characteristics collectively ensure that the selected datasets offer a representative and challenging environment for comparing the performance and computational efficiency of the proposed Irredundant $k$–fold validation against the standard $k$–fold approach. A concise overview of these datasets appers in Table 1.
All experiments used 5 folds ($k\!=\!5$), and employed Random Forest classifier with default hyper–parameters implemented in scikit--learn Python package [28], and were run on an Apple M1 Ultra (20 cores).
Table 1 presents a detailed comparison between the Irredundant $k$–fold ($\mathsf{IkF}$) and the standard $k$–fold ($\mathsf{kF}$) cross–validation methods across a diverse selection of datasets. For each method, we report the obtained accuracy (Acc), F–score (Fsc), and computational Time (in seconds). Additionally, we provide the relative ratios $\mathsf{IkF}/\mathsf{kF}$ for each metric, namely RAcc, RFsc, and Speed–up. The last row shows the average values across all datasets.
[…]
The consistency of these findings over datasets that vary widely in sample size, dimensionality, and class cardinality underlines the robustness of $\mathsf{IkF}$ Consequently, the irredundant method offers a reliable, efficient alternative to classical $\mathsf{kF}$—particularly valuable when resources are constrained or rapid model assessment is required.

### v2

@@ Abstract
In traditional k–fold cross–validation, each instance is used (k1k\!-\!1) times for training and once for testing, leading to redundancy that lets many instances disproportionately influence the learning phase. We introduce Irredundant kk–fold cross–validation, a novel method that guarantees each instance is used exactly once for training and once for testing across the entire validation procedure. This approach ensures a more balanced utilization of the dataset, mitigates overfitting due to instance repetition, and enables sharper distinctions in comparative model analysis. The method preserves stratification and remains model–agnostic, i.e., compatible with any classifier. Experimental results demonstrate that it delivers consistent performance estimates across diverse datasets –comparable to kk–fold cross–validation– while providing less optimistic variance estimates because training partitions are non–overlapping, and significantly reducing the overall computational cost.
@@ 1 Introduction
The hold–out method [7] improves on this by partitioning the data into two disjoint subsets: one for training and another for testing (typically, $2/3$ for training and $1/3$ for testing). The hold–out error estimator is unbiased for the generalization error provided that the test data are truly independent of the training data and have not been reused in any form [8, 9]. While computationally inexpensive, the hold–out method is sensitive to how the data are split, resulting in high variance in the estimated metrics.
To address the instability of hold–out, kk–fold cross–validation was proposed [10]. In this method, the data is partitioned into $k$ disjoint subsets (folds). Each fold serves once as a test set, while the remaining k1k-1 folds are used for training. The final performance estimate is the average across all $k$ iterations. Cross–validation reduces variance and produces more robust estimates, particularly in small datasets. However, its variance is significantly greater than that of resubstitution [11]. As noted by Nadeau & Bengio [12], the use of cross–validation estimators for model selection sparked a debate [13, 14] related to the “no free lunch theorem” [15], highlighting that while cross–validation often works well in practice, it is not guaranteed to perform uniformly well across all learning problems.
An important extension of kk–fold cross–validation is nested cross–validation, in which an inner cross–validation loop (usually (k1)(k\!-\!1)–fold) is executed within each outer test fold, and the risk estimate is adjusted using the outer–fold loss [16]. This recent approach provides a more reliable estimate of generalization error by accounting for the additional variability introduced by model selection or hyperparameter tuning.
A limiting case of kk–fold cross–validation is leave–one–out cross–validation [17, 18, 19], where each individual instance serves once as the test instance while the rest are used for training. Although this method yields minimal bias, it can be computationally intensive and prone to high variance, particularly for models with high complexity.
A generalization of leave–one–out is leave–pp–out cross–validation [20], where $p$ instances are left out at each iteration, and every possible subset of $p$ instances is used as a validation set. Notably, leave–one–out corresponds to leave–pp–out with p=1p\!=\!1.
@@
Each validation strategy entails trade–offs in terms of computational cost, variance, and bias. Hold–out is fast but unstable; leave–one–out is unbiased but often computationally prohibitive; kk–fold cross–validation balances variance and bias reasonably well but can still be expensive; and bootstrap methods introduce an additional stochastic layer, sometimes at the cost of increased bias, that may not always reflect real–world generalization.
As models increase in complexity and datasets become larger and more heterogeneous, more sophisticated and computationally efficient validation schemes are needed. This paper introduces irredundant kk–fold cross–validation, a novel approach that aims to combine statistical rigor with computational efficiency. Unlike standard kk–fold cross–validation, where each sample is used k1k\!-\!1 times for training, irredundant kk–fold ensures that each sample is used exactly once for training and exactly once for testing. Experimental results demonstrate that irredundant kk–fold achieves significant computational savings while preserving the quality of performance estimates.
@@
The remainder of this paper is organized as follows: Section 2 describes the irredundant kk–fold cross–validation in detail and includes, in Section 3, a brief analysis of its bias and variance; Section 4 presents empirical results across several dataset from the UCI repository. Finally, 5 summarizes the main conclusions and outlines future work.
@@ 2.1 Overview
The fundamental difference between standard kk–fold cross–validation ($\mathsf{kF}$) and the proposed irredundant kk–fold cross–validation ($\mathsf{IkF}$) lies how many times each instance is used for training. In traditional $\mathsf{kF}$, each instance appears in the training set (k1k-1) out of the $k$ iterations. Therefore, per iteration:
@@
To achieve this, $\mathsf{IkF}$ introduces a subfolding strategy. Each outer fold is divided into k1k-1 subfolds. For each test fold $\mathcal{F}^{i}$, the training set $\mathcal{E}^{i}$ is constructed by selecting one unique, unused subfold from each fold $\mathcal{F}_{j}$ with $j\neq i$. The subfolds may be selected randomly or via heuristic criteria, as long as each is used exactly once across all folds.
@@
[…]
@@
The proposed $\mathsf{IkF}$ guarantees that each instance appears exactly once in a test set $\mathcal{T}^{i}$ and exactly once in a training set $\mathcal{E}^{i}$. Furthermore, each fold $\mathcal{F}_{i}$ is used exactly once for testing, and each training set $\mathcal{E}^{i}$ is composed of k1k\!-\!1 distinct subfolds, one from each of the other folds. Thus, the size of each training set remains approximately $n/k$.
@@ 2.2 Algorithm Description
The full $\mathsf{IkF}$ procedure is summarized in Algorithm 1. Structurally, it resembles classical $\mathsf{kF}$, but differs in Line #7, where the training set $\mathcal{E}^{i}$ is built by assembling one unused subfold from each of the k1k-1 non–test folds.
@@
3:  Split $\mathcal{F}_{i}$ into kk-11 subfolds $\mathcal{F}_{i}^{j}$: $\mathcal{F}_{i}=\bigcup_{j=1}^{k-1}\mathcal{F}_{i}^{j}$
@@
6:  $\mathcal{T}^{i}\!\leftarrow\!\mathcal{F}_{i}$ $\triangleright$ Test set
7:  $\mathcal{E}^{i}\!\leftarrow\!\bigcup_{j\neq i}\mathcal{F}_{j}^{\alpha_{ij}}$ where $\alpha_{ij}\!\in\!\{1,\!\dots\!,k\!-\!1\}$ $\triangleright$ Training set
@@
9:  $\theta^{i}$ $\leftarrow$ Evaluate $\mathcal{M}^{i}$ on $\mathcal{T}^{i}$ with $\delta$ $\triangleright$ Performance
@@
Because each fold is split into k1k\!-\!1 subfolds, the expected size of a subfold is n/[k(k1)]n/[k(k-1)]. Hence the irredundant scheme is feasible only if $n\geq k(k-1)$, which is easily met by most real–world data.
@@
Figure 1: Graphical comparison between kk–fold cross–validation and Irredundant kk–fold cross–validation.
Figure 1 illustrates the full $\mathsf{IkF}$ construction for $k=3$. To the left, the traditional kk–fold cross–validation; in the center, the decomposition for irredundant kk–fold cross–validation into $k$ folds, which of them is again decomposed into k1k\!-\!1 subfolds.
@@ 2.3 Computational cost
A core advantage of $\mathsf{IkF}$ is its lower computational cost. Since each training set is smaller (size $n/k$ vs. (k1)n/k(k\!-\!1)n/k), and each instance is used for training only once, overall computational demand is reduced.
@@
The relative cost reduction depends on the model’s complexity. For linear–time learners (e.g., naïve Bayes), $\mathcal{C}_{\mathsf{kF}}/\mathcal{C}_{\mathsf{IkF}}\!=\!k\!-\!1$. In a typical 10–fold setup, this means up to $9\times$ speed–up with respect to the classical $\mathsf{kF}$ approach.
@@ 3 Analysis
In a generic kk–fold cross-validation scheme—either standard or the proposed irredundant—the dataset $\mathcal{D}$ is partitioned into $k$ mutually–exclusive training–test pairs $(\mathcal{E}_{1},\mathcal{T}_{1}),\;(\mathcal{E}_{2},\mathcal{T}_{2}),\;\dots,\;(\mathcal{E}_{k},\mathcal{T}_{k})$, where each $\mathcal{T}_{i}$ serves as the validation set for the model trained on $\mathcal{E}_{i}$. The fold–wise error estimate and the resulting cross–validation estimator are defined as follows:
@@
[…]
Under mild regularity assumptions, both standard and irredundant kk–fold cross–validation yield consistent estimators of the generalization risk. Let $R(h_{n})$ denote the risk of a model trained on $n$ i.i.d. samples. If the learning algorithm $\mathcal{M}$ is universally consistent, i.e., $R(h_{n})\to R^{*}$ as $n\to\infty$ for all $\mathbb{P}_{XY}$, and the loss function $\ell$ is bounded, then the cross–validation estimate $\widehat{\theta}$ converges in probability to $R^{*}$. In particular, for the proposed irredundant procedure $\mathsf{IkF}$, the training set size per fold is $n/k$, which grows with $n$ when $k$ is fixed. Therefore, $\mathsf{IkF}$ remains consistent, since each model in the ensemble is trained on increasingly large samples. Moreover, the absence of training–set overlap in $\mathsf{IkF}$ reduces inter–fold dependence, simplifying variance analysis and enhancing reliability in finite samples. Consequently, $\mathsf{IkF}$ inherits the desirable asymptotic properties of classical cross–validation while providing practical advantages in terms of computational efficiency and variance reduction.
@@ 3.1 Bias Analysis
[…]
@@ 3.2 Variance Analysis
[…]
High variance indicates that the model’s performance estimate is highly sensitive to the chosen data partition, which undermines the robustness of conclusions. In traditional $\mathsf{kF}$ cross–validation, each instance is used for training in k1k\!-\!1 folds and for testing in one fold. Because training sets might be large and largely overlapping, the variance between fold estimates is often low—primarily due to the shared training samples across iterations. However, this overlap reduces the diversity of the training sets and may lead to over–optimistic estimates of generalization.
In general, the use of standard $\mathsf{kF}$ can lead to uneven model behavior, mainly due to redundant learning: since each instance appears in the training set of k1k\!-\!1 folds, the same examples are learned repeatedly. This redundancy artificially lowers the observed variance in performance estimates, as the model benefits multiple times from having seen the same samples. As a result, $\mathsf{kF}$ can underestimate the variance that would be observed in real deployment on truly unseen data [26].
@@
However, this increased per–fold variance is counterbalanced by the elimination of inter–fold covariance. In standard $\mathsf{kF}$, the proportion of overlap between any two training sets is exactly (k2)/k(k-2)/k; for example, 80% when $k=10$. This overlap induces a positive covariance between fold estimates, thereby inflating the overall variance of the estimator. In $\mathsf{IkF}$, the training sets are completely disjoint, and the test sets do not influence other iterations, so the covariances $\text{Cov}(\hat{\theta}_{i},\hat{\theta}_{j})$ are negligible or zero.
@@ 3.3 Bias–Variance Trade–off
[…]
@@
Generally, the variance of $\hat{\theta}$ tends to increase with the number of folds kk—since each test set becomes smaller and each training set slightly more variable—and decrease as $k$ decreases (i.e., when test sets are larger and folds are more stable). At small $k$ (e.g., $k=5$), bias increases because each model is trained on significantly less data, but variance decreases due to lower instability. At the other extreme, leave-one-out cross-validation ($k=N$) yields minimal bias—since each model is trained on nearly all data—but high variance due to high sensitivity to the specific omitted sample.
@@ 4 Experimental Analysis
Table 1: Performance of the Irredundant kk–fold CV ($\mathsf{IkF}$) and the standard kk–fold CV ($\mathsf{kF}$) methods across multiple datasets using the Random Forest classifier. #s, #v and #c stand for the number of samples, variables and classes, respectively. For each method, accuracy (Acc), F–score (Fsc), and running Time (in sec.) are reported. Relative ratios $\mathsf{kF}$$\mathsf{IkF}$ are shown in the last three columns. Last row shows the mean of each column.
Irredundant kk–Fold Standard kk–Fold Ratios $\mathsf{kF}/\mathsf{IkF}$
@@
We benchmarked $\mathsf{IkF}$ and $\mathsf{kF}$ on ten publicly available datasets from the UCI Machine–Learning Repository [27]. Several properties make these datasets particularly suitable for the present analysis. First, they vary substantially in sample size, from relatively small datasets (e.g., hepatitis_c with 1,385 samples) to larger datasets (e.g., magic_gamma with 19,020 samples), allowing the evaluation of validation methods across different sizes. Second, the number of predictive variables covers a wide spectrum, from low–dimensional problems (e.g., data_banknote with 4 variables) to high–dimensional ones (e.g., musk_v2 with 166 variables), providing a testbed for studying the impact of dimensionality. Third, the classification tasks include both binary and multiclass problems, ensuring that the methods are assessed under varying levels of label complexity. These characteristics collectively ensure that the selected datasets offer a representative and challenging environment for comparing the performance and computational efficiency of the proposed Irredundant kk–fold validation against the standard kk–fold approach. A concise overview of these datasets appers in Table 1.
All experiments used 5 folds (k=5k\!=\!5), and employed Random Forest classifier with default hyper–parameters implemented in scikit-learn Python package [28], and were run on an Apple M1 Ultra (20 cores).
Table 1 presents a detailed comparison between the Irredundant kk–fold ($\mathsf{IkF}$) and the standard kk–fold ($\mathsf{kF}$) cross–validation methods across a diverse selection of datasets. For each method, we report the obtained accuracy (Acc), F–score (Fsc), and computational Time (in seconds). Additionally, we provide the relative ratios $\mathsf{kF}/\mathsf{IkF}$ for each metric, namely RAcc, RFsc, and Speed–up. The last row shows the average values across all datasets.
The results clearly demonstrate that the $\mathsf{IkF}$ approach achieves performance metrics (accuracy and F–score) that are closely comparable to those obtained by the standard $\mathsf{kF}$ validation. The mean accuracy ratio $\text{RAcc}=1.027$ and the mean F–score ratio $\text{RFsc}=1.048$ are both close to 1, indicating that, on average, removing redundant training instances does not degrade performance.
The performance was also similar when probabilistic measures were used. For instance, for the multiclass classification performance (MCP) curve [29, 30, 31], the area under the curve was calculated, providing a mean value of 0.751 for $\mathsf{IkF}$ and of 0.792 for $\mathsf{kF}$, which is consistent with the two previous ones.
A key advantage of $\mathsf{IkF}$ lies in its significant reduction in computational cost. On average, it reduces the total validation time by a factor of approximately 2.9 compared to $\mathsf{kF}$, as reflected by the mean Speed–up value. Equivalent gains were also observed (results not shown) with two additional learners—kk–nearest neighbours and Naïve–Bayes—indicating that the savings are classifier–agnostic. Crucially, this acceleration is obtained without meaningful loss of predictive quality: the average accuracy and F–score differ from those of standard $\mathsf{kF}$ by only +2.7% and +4.8%, respectively.
The consistency of these findings over datasets that vary widely in sample size, dimensionality, and class cardinality underlines the robustness of $\mathsf{IkF}$ . Consequently, the irredundant method offers a reliable, efficient alternative to classical $\mathsf{kF}$—particularly valuable when resources are constrained or rapid model assessment is required.